a bullet fired from a gun was given by f equals to 600 - 2 into 10 ^ fees in Newtons and 30 seconds the falls on the budget becomes zero as soon as the Bala what is the average impulse imparted to the bullet
 

Grade:9

2 Answers

Arun
4 years ago
F = 600 - 2 * 10⁵t
0 = 600 - 2 * 10⁵ t
t = 600 / (2 * 10⁵)
t = 3 * 10⁻³ s
t = 3 milliseconds
Time taken by bullet to leave the barrel is 3 milliseconds
F.dt = (600 - 2* 10⁵t).dt
J = [600t - (2 * 10⁵ t²/2)]
J = [600*(3*10^-3) - (2 * 10^5 * (3 * 10^-3)^2 / 2)]
J = 0.9 Ns
Impulse imparted to bullet is 0.9 Ns
 
Vikas TU
4 years ago
Dear student 
Above ans is incorrect 
The force will become zero in amount of time 't'
so,
F =   600 - 200000t = 0
or
t = 600 / 200000
thus,
t = 3 x 10 -3 s
now, we know that
Impulse = ∫Fdt = ∫(   600 - 200000t ).dt
which gives us
I = 1.8 Ns
Hope this helps